# A Novel Garment Transfer Method Supervised by Distilled Knowledge of Virtual Try-on Model

### v1

@@ 4.1 Implementation Details
We exploit the Zalando dataset (Han et al., 2018) to train and evaluate our proposed method [^2] We have gotten the using permission for the research purpose from its author Xintong Han.. This dataset is further cleaned so that the training set and test set consist of 11565 and 1698 paired data with a resolution of ${256\times 192}$. We randomly match person ${A}$ and ${B}$ to form the garment transfer pair. All experiments were conducted on a single 3090 NVIDIA GPU by Adam optimizer. We present the hyperparameters of garment transfer models in the training phase in Table 1 and provide the model complexity in Table 2.
@@ 4.2.1 Transfer Parsing Reasoning
Metrics In this section, we conduct the ablation study of transfer parsing reasoning to verify the necessity of knowledge distillation in training ${{{\cal S}_{p}}}$, the rationality of the phase setting in the teaching procedure, and the appropriateness of the backbone and pose form for this task. We exploit the intersection over union (IoU) and class pixel accuracy (CPA) as metrics, both of which range from 0 to 1. Since the entanglement between the garment and pose is crucial in transfer parsing reasoning, we focus on the metrics values of the garment and upper body skin categories. Thus, we compare ${{{\cal M}_{SAB}}}$ with ${{{\cal M}_{TAB}}}$ to verify the knowledge distillation effect, denoted as ${{\rm{IoU}}\left({{\cal M}_{TAB}^{ha},{\cal M}_{SAB}^{ha}}\right)}$,${{\rm{IoU}}\left({{\cal M}_{TAB}^{hc},{\cal M}_{SAB}^{hc}}\right)}$, ${{\rm{CPA}}\left({{\cal M}_{TAB}^{ha},{\cal M}_{SAB}^{ha}}\right)}$, and ${{\rm{CPA}}\left({{\cal M}_{TAB}^{hc},{\cal M}_{SAB}^{hc}}\right)}$. And we compare ${{{\cal M}_{SB}}}$ with ${{{\cal M}_{B}}}$ to verify the reasoning rationality with ground truth, denoted as ${{\rm{IoU}}\left({{\cal M}_{SB}^{ha},{\cal M}_{B}^{ha}}\right)}$,${{\rm{IoU}}\left({{\cal M}_{SB}^{hc},{\cal M}_{B}^{hc}}\right)}$, ${{\rm{CPA}}\left({{\cal M}_{SB}^{ha},{\cal M}_{B}^{ha}}\right)}$, and ${{\rm{CPA}}\left({{\cal M}_{SB}^{hc},{\cal M}_{B}^{hc}}\right)}$.
The Necessity of Knowledge Distillation As described in Sec. 3.2.1, without knowledge distillation, we train ${{{\cal S}_{p}}}$ in a self-supervised manner as comparisons. Specifically, during the training phase, ${{{\cal S}_{p}}}$ is conditioned on augmented ${{\cal M}_{A}^{hc}}$, ${{\cal M}_{A}^{hr}}$, and ${{{\cal P}_{A}}}$, while the ground truth ${{{\cal M}_{A}}}$ is leveraged to supervise the output of ${{{\cal S}_{p}}}$. The traditional data augmentation operations, including affine transformation, cropping, and flipping, are used in three different combinations: affine transformation; affine transformation and cropping; and affine transformation, cropping, and flipping. To further improve the performance, we adopted the cycle consistency strategy from (Ge et al., 2021a), which involves dressing ${{\cal M}_{B}^{hc}}$ onto person ${A}$ to generate ~${{\tilde{\cal M}_{AB}}}$ and then dressing ~${\tilde{\cal M}_{AB}^{hc}}$ onto person ${B}$ again, supervised by the ground truth ${{{\cal M}_{B}}}$. The quantitative and qualitative results are shown in Fig. 8 and Fig. 9, respectively.
@@
The Appropriateness of Backbone and Pose Form We compare the appropriateness of different backbones for the parsing reasoning task, including U2-Net (Qin et al., 2020), U-Net (Ronneberger et al., 2015), ResUnet (Zhang et al., 2018), and UNet++ (Zhou et al., 2019). We compare the function of DensePose and OpenPose in reasoning. The quantitative and qualitative results are shown in Fig. 12 and Fig. 13, respectively.
@@
The U2-Net backbone, which utilizes a deep architecture by nesting U-Net, is adept at capturing semantic features of garments and poses. It excels in reconstructing and propagating local details by fusing multiscale features, resulting in better preservation of arm shape and garment details compared to other backbones, as indicated by a reduction of ${{\rm{IoU}}\left({{\cal M}_{SB}^{hc},{\cal M}_{B}^{hc}}\right)}$ and ${{\rm{IoU}}\left({{\cal M}_{SB}^{hc},{\cal M}_{B}^{hc}}\right)}$ by 6.87% and 2.60%, respectively. Thus, U2-Net is deemed more suitable for the parsing reasoning task among the considered backbones. On the other hand, OpenPose only provides information on joint positions, and thus, fails to reconstruct fine arm shapes due to the lack of arm shape prior. In contrast, DensePose, which provides human structure and shape prior, results in an increase of ${{\rm{IoU}}\left({{\cal M}_{SAB}^{ha},{\cal M}_{TAB}^{ha}}\right)}$ and ${{\rm{IoU}}\left({{\cal M}_{B}^{ha},{\cal M}_{SB}^{ha}}\right)}$ by 5.51% and 3.34%, respectively.
@@ 4.2.2 Garment Warping
Metrics In this section, we verify the necessity of knowledge distillation in training , the appropriateness of warping method, and the rationality of phase setting. Similar to Sec. 4.2.1, we compare ~${\tilde{\cal I}_{SAB}^{hc}}$ with ~${\tilde{\cal I}_{TAB}^{hc}}$ to verify the knowledge distillation and compare ~${\tilde{\cal I}_{SB}^{hc}}$ with ${{\cal I}_{B}^{hc}}$ to verify the garment warping authenticity to the ground truth. We exploit the low-level full-reference metrics, high-level full-reference metrics, and no-reference subjective metric. The low-level full-reference metrics consist of MSE, SSIM, and PSNR; the high-level full-reference metrics consist of FID (Heusel et al., 2017) and LPIPS (Zhang et al., 2018); the no-reference subjective metric is IS (Salimans et al., 2016).
@@
The Appropriateness of Warping Method As described in Sec. 3.3.1, we evaluate the appropriateness of mapping, TPS deformation, and TPS deformation+render for garment transfer. To ensure a fair comparison, we employ the same knowledge distillation pipeline when training these methods. Specifically, for the mapping method, we leverage ${{{\cal T}_{w}}}$ as the teacher model and the texture module of SwapNet (Raj et al., 2018) as the student model. The student model is conditioned on ${{\cal M}_{SAB}^{hc}}$ and the garment embedding, which is obtained by ROI pooling of ${{\cal I}_{B}^{hc}}$; In the case of TPS deformation, we exploit the clothes warping module of ACGPN (Yang et al., 2020) as the teacher and student models. The student model is conditioned on ${{\cal M}_{B}^{hc}}$ and ${{\cal M}_{SAB}^{hc}}$. In the case of TPS deformation + render, we further input the warped result of case III into a refinement network, which is the same as ACGPN. The quantitative and qualitative results are shown in Fig. 16 and 17.
@@
Case II is burdened with complex texture mapping due to variations in the texture and decoration of the garment. Consequently, an average color is padded into the target shape, resulting in blurred texture and discarded decoration. This causes an increase of 85.54% and 92.51% in ${{\rm{FID}}\left({\tilde{\cal I}_{TAB}^{hc},\tilde{\cal I}_{SAB}^{hc}}\right)}$ and ${{\rm{FID}}\left({\tilde{\cal I}_{SB}^{hc},{\cal I}_{B}^{hc}}\right)}$, respectively. As TPS deformation has a low DOF, while TPS deformation works well in obtaining a coarse warped result for garments with standard poses and large scales in virtual try-on, fine warping from pose to pose is necessary for garment transfer. Therefore, case III cannot accurately warp the garment to match the target shape and tends to produce distortion, leading to an increase of 175.44% and 341.51% in ${{\rm{MSE}}\left({\tilde{\cal I}_{TAB}^{hc},\tilde{\cal I}_{SAB}^{hc}}\right)}$ and ${{\rm{FID}}\left({\tilde{\cal I}_{TAB}^{hc},\tilde{\cal I}_{SAB}^{hc}}\right)}$, respectively. In case IV, rendering texture is also a challenge, and the poor warped result and unaligned ${{\cal I}_{B}^{hc}}$ cause training instability and collapse, resulting in a reduction of ${{\rm{IS}}\left({\tilde{\cal I}_{SAB}^{hc}}\right)}$ by 24.01%.
@@
In case II, position mapping and shape adjustment are accomplished through high-DOF flow warping. However, when ${{\cal M}_{B}^{hc}}$ is significantly smaller than ${{\cal M}_{SAB}^{hc}}$, it tends to excessively stretch the local region to match the target shape precisely. In contrast, case I aligns ${{\cal I}_{B}^{hc}}$ globally with ${{\cal M}_{SAB}^{hc}}$ to achieve a similar scale and position. Subsequently, flow warping is leveraged to adjust the shape accurately. As a result, the texture blur and garment style variation are reduced, and ${{\rm{MSE}}\left({\tilde{\cal I}_{TAB}^{hc},\tilde{\cal I}_{SAB}^{hc}}\right)}$ and ${{\rm{LPIPS}}\left({\tilde{\cal I}_{TAB}^{hc},\tilde{\cal I}_{SAB}^{hc}}\right)}$ decrease by 4.32% and 2.27%. In case III, the FSV modulates the pose and garment features to estimate a coarse try-on result, which performs well in virtual try-on without parsing reasoning. However, our garment transfer pipeline provides the fined target shape, rendering the mechanism of FSV unnecessary. Conversely, the innovation of case I learns correspondence at the shape and content level, emphasizing semantic alignment and consistency in garment warping. As a result, ~${\tilde{\cal I}_{SAB}^{hc}}$ is more realistic, and ${{\rm{IS}}\left({\tilde{\cal I}_{SAB}^{hc}}\right)}$ and ${{\rm{IS}}\left({\tilde{\cal I}_{SB}^{hc}}\right)}$ increase by 2.05% and 3.15%.
@@ 4.2.3 Arm Regrowth
The task of arm regrowth can be divided into two sub-tasks: propagating the original arm part and reasoning the new arm part. In case II, the augmented ${{\cal I}_{A}^{ha}}$ is misaligned with ${{\cal M}_{A}^{ha}}$ through an affine transformation, resulting in ${{G_{r}}}$ padding the overall target shape with the semantic features of ${{\cal I}_{A}^{ha}}$. Consequently, the results of case II discard high-frequency details such as the hand and neck by inferring rather than propagating the content at the overlapped region between ${{\cal M}_{SAB}^{ha}}$ and ${{\cal M}_{A}^{ha}}$. In case III, new arm exposure is more realistically simulated in the training phase by using random cropping, resulting in a relative reduction of 18.85% and 14.54% in MSE and LPIPS respectively when compared to case II. Nevertheless, the results of case III still exhibit artifacts and blurring. Case IV further impedes the propagation of the original arm part by using random flipping, resulting in worse results compared to case III. Unlike these data augmentation techniques, case I aims to enable the input conditions of the training and testing phases as similar as possible to perform well in both sub-tasks, leading to significant quantitative results.
@@ 4.3.1 Virtual Try-on Method
Baseline We compare our method with virtual try-on methods in garment transfer. Virtual try-on methods consist of VITON (Han et al., 2018), CP-VTON (Wang et al., 2018), ACGPN (Yang et al., 2020), DCTON (Ge et al., 2021a), PF-AFN (Ge et al., 2021b), and FSV (He et al., 2022). We adopt their official implementations and checkpoint pre-trained on the Zalando dataset [^3] Since CP-VTON does not release its official code, we adopt the most starred unofficial code on Github (https://github.com/sergeywong/cp-vton).
Metrics We compare the performance of different methods for transferring the garment from ${{{\cal I}_{B}}}$ to ${{{\cal I}_{A}}}$ to obtain ~${{\tilde{\cal I}_{AB}}}$. In order to perform this transfer, virtual try-on methods are conditioned on both ${{{\cal I}_{A}}}$ and ${{\cal I}_{B}^{hc}}$. As the ground truth for ~${{\tilde{\cal I}_{AB}}}$ is unavailable, we present several examples in Fig. 21 and evaluate their subjective image quality using IS (Salimans et al., 2016) and hyperIQA (Su et al., 2020) in Table 4. To conduct a more detailed quantitative comparison, we re-transfer the garment from ~${{\tilde{\cal I}_{AB}}}$ to ${{{\cal I}_{B}}}$ and utilize ${{{\cal I}_{B}}}$ as the ground truth. We also adopt the same metrics as in Sec. 4.2.2. Specifically, we utilize the pre-trained Grapy-ML to extract ~${{\tilde{\cal M}_{AB}}}$ from ~${{\tilde{\cal I}_{AB}}}$ and yield the transferred garment ~${\tilde{\cal I}_{AB}^{hc}}$. Subsequently, virtual try-on methods are conditioned on both ~${\tilde{\cal I}_{AB}^{hc}}$ and ${{{\cal I}_{B}}}$ to predict ~${{\tilde{\cal I}_{B}}}$, and the quantitative results are presented in Table 5.
@@
Table 4: Quantitative comparison with virtual try-on methods in garment transfer from ${{{\cal I}_{B}}}$ to ${{{\cal I}_{A}}}$.
Ours VITON CP-VTON ACGPN DCTON PF-AFN FSV
IS↑ 2.75 2.39 2.42 2.50 2.42 2.31 2.38
hyperIQA↑ 35.94 31.81 28.20 27.11 29.49 28.80 27.92
Table 5: Quantitative comparison with virtual try-on methods in garment transfer from ~${{\tilde{\cal I}_{AB}}}$ to ${{{\cal I}_{B}}}$.
MSE↓ PSNR↑ SSIM↑ FID↓ LPIPS↓ IS↑
Ours 389.029 23.57 0.863 19.14 0.108 2.734
VITON 3867.18 16.15 0.700 75.91 0.258 2.374
CP-VTON 1853.72 16.27 0.712 53.10 0.244 2.266
ACGPN 647.72 21.31 0.825 27.64 0.147 2.496
DCTON 770.71 20.27 0.814 32.18 0.159 2.455
PF-AFN 1359.94 18.15 0.823 46.15 0.176 2.280
FSW 1711.46 17.12 0.820 56.50 0.191 2.317
Virtual try-on methods are limited in their ability to capture position and style features of transferred garments, resulting in warped garments that do not align with the transfer parsing and yielding artifacts in the unaligned region. In contrast, our method precisely warps ${{\cal I}_{B}^{hc}}$ towards ${{\cal M}_{SAB}^{hc}}$ through initial aligning and flow warping. Additionally, these virtual try-on methods cannot avoid being affected by the garment shape features of ${{{\cal I}_{A}}}$ when reasoning a proper transfer parsing, leading to ${{\cal I}_{B}^{hc}}$ warping toward ${{\cal I}_{A}^{hc}}$. In contrast, our method infers a parsing that entangles the shape feature of the transfer garment with the pose feature of another person, while maintaining their inherent features. This allows our results to be more aligned with the intended garment transfer. In quantitative comparisons, our method achieves the highest scores of IS and hyperIQA in garment transfer from ${{{\cal I}_{B}}}$ to ${{{\cal I}_{A}}}$. Moreover, our MSE and FID are approximately 20.89% and 34.37% lower than those of other virtual try-on methods in garment transfer from ~${{\tilde{\cal I}_{AB}}}$ to ${{{\cal I}_{B}}}$.

### v2

@@ 4.1 Implementation Details
We exploit the Zalando dataset [1] to train and evaluate our proposed method [^2] We have gotten the using permission for the research purpose from its author Xintong Han.. This dataset is further cleaned so that the training set and test set consist of 11565 and 1698 paired data with a resolution of ${256\times 192}$. We randomly match person ${A}$ and ${B}$ to form the garment transfer pair. All experiments were conducted on a single 3090 NVIDIA GPU by Adam optimizer. We present the hyperparameters of garment transfer models in the training phase in Table 1 and provide the model complexity in Table 2.
@@ 4.2.1 Transfer Parsing Reasoning
Metrics In this section, we conduct the ablation study of transfer parsing reasoning to verify the necessity of knowledge distillation in training ${{{\cal S}_{p}}}$, the rationality of the phase setting in the teaching procedure, and the appropriateness of the backbone and pose form. We exploit the intersection over union (IoU) and class pixel accuracy (CPA) as metrics, both of which range from 0 to 1. Since the entanglement between the garment and pose is crucial in transfer parsing reasoning, we focus on the metrics values of the garment and upper body skin categories. Thus, we compare ${{{\cal M}_{SAB}}}$ with ${{{\cal M}_{TAB}}}$ to verify the knowledge distillation effect, denoted as ${{\rm{IoU}}\left({{\cal M}_{TAB}^{ha},{\cal M}_{SAB}^{ha}}\right)}$, ${{\rm{IoU}}\left({{\cal M}_{TAB}^{hc},{\cal M}_{SAB}^{hc}}\right)}$, ${{\rm{CPA}}\left({{\cal M}_{TAB}^{ha},{\cal M}_{SAB}^{ha}}\right)}$, and ${{\rm{CPA}}\left({{\cal M}_{TAB}^{hc},{\cal M}_{SAB}^{hc}}\right)}$. And we compare ${{{\cal M}_{SB}}}$ with ${{{\cal M}_{B}}}$ to verify the reasoning rationality with ground truth, denoted as ${{\rm{IoU}}\left({{\cal M}_{SB}^{ha},{\cal M}_{B}^{ha}}\right)}$, ${{\rm{IoU}}\left({{\cal M}_{SB}^{hc},{\cal M}_{B}^{hc}}\right)}$, ${{\rm{CPA}}\left({{\cal M}_{SB}^{ha},{\cal M}_{B}^{ha}}\right)}$, and ${{\rm{CPA}}\left({{\cal M}_{SB}^{hc},{\cal M}_{B}^{hc}}\right)}$.
The Necessity of Knowledge Distillation As described in Sec. 3.2.1, without knowledge distillation, we train ${{{\cal S}_{p}}}$ in a self-supervised manner as comparisons. Specifically, during the training phase, ${{{\cal S}_{p}}}$ is conditioned on augmented ${{\cal M}_{A}^{hc}}$, ${{\cal M}_{A}^{hr}}$, and ${{{\cal P}_{A}}}$, while the ground truth ${{{\cal M}_{A}}}$ is leveraged to supervise the output of ${{{\cal S}_{p}}}$. The traditional data augmentation operations, including affine transformation, cropping, and flipping, are used in three different combinations: affine transformation; affine transformation and cropping; and affine transformation, cropping, and flipping. To further improve the performance, we adopted the cycle consistency strategy from [8], which involves dressing ${{\cal M}_{B}^{hc}}$ onto person ${A}$ to generate ~${{\tilde{\cal M}_{AB}}}$ and then dressing ~${\tilde{\cal M}_{AB}^{hc}}$ onto person ${B}$ again, supervised by the ground truth ${{{\cal M}_{B}}}$. The quantitative and qualitative results are shown in Fig. 8 and Fig. 9, respectively.
@@
The Appropriateness of Backbone and Pose Form We compare the appropriateness of different backbones for the parsing reasoning task, including U2-Net [30], U-Net [31], ResUnet [29], and UNet++ [32]. We compare the function of DensePose and OpenPose in reasoning. The quantitative and qualitative results are shown in Fig. 12 and Fig. 13, respectively.
@@
The U2-Net backbone utilizes a deep architecture by nesting U-Net and is adept at capturing semantic features of garments and poses. It excels in reconstructing and propagating local details by fusing multiscale features, resulting in better preservation of arm shape and garment details compared to other backbones, as indicated by a reduction of ${{\rm{IoU}}\left({{\cal M}_{SB}^{hc},{\cal M}_{B}^{hc}}\right)}$ and ${{\rm{IoU}}\left({{\cal M}_{SB}^{hc},{\cal M}_{B}^{hc}}\right)}$ by 6.87% and 2.60%, respectively. Thus, U2-Net is deemed more suitable for the parsing reasoning task among the considered backbones. On the other hand, OpenPose only provides information on joint positions, and thus, fails to reconstruct fine arm shapes due to the lack of arm shape prior. In contrast, DensePose provides human structure and shape prior and results in an increase of ${{\rm{IoU}}\left({{\cal M}_{SAB}^{ha},{\cal M}_{TAB}^{ha}}\right)}$ and ${{\rm{IoU}}\left({{\cal M}_{B}^{ha},{\cal M}_{SB}^{ha}}\right)}$ by 5.51% and 3.34%, respectively.
@@ 4.2.2 Garment Warping
Metrics In this section, we verify the necessity of knowledge distillation in training , the appropriateness of warping method, and the rationality of phase setting. Similar to Sec. 4.2.1, we compare ~${\tilde{\cal I}_{SAB}^{hc}}$ with ~${\tilde{\cal I}_{TAB}^{hc}}$ to verify the knowledge distillation and compare ~${\tilde{\cal I}_{SB}^{hc}}$ with ${{\cal I}_{B}^{hc}}$ to verify the garment warping authenticity to the ground truth. We exploit the low-level full-reference metrics, high-level full-reference metrics, and no-reference subjective metric. The low-level full-reference metrics consist of MSE, SSIM, and PSNR; the high-level full-reference metrics consist of FID [33] and LPIPS [34]; the no-reference subjective metric is IS [35].
@@
The Appropriateness of Warping Method As described in Sec. 3.3.1, we evaluate the appropriateness of mapping, TPS deformation, and TPS deformation+render for garment transfer. To ensure a fair comparison, we employ the same knowledge distillation pipeline when training these methods. Specifically, for the mapping method, we leverage ${{{\cal T}_{w}}}$ as the teacher model and the texture module of SwapNet [3] as the student model. The student model is conditioned on ${{\cal M}_{SAB}^{hc}}$ and the garment embedding, which is obtained by ROI pooling of ${{\cal I}_{B}^{hc}}$; In the case of TPS deformation, we exploit the clothes warping module of ACGPN [5] as the teacher and student models. The student model is conditioned on ${{\cal M}_{B}^{hc}}$ and ${{\cal M}_{SAB}^{hc}}$. In the case of TPS deformation + render, we further input the warped result of case III into a refinement network, which is the same as ACGPN. The quantitative and qualitative results are shown in Fig. 16 and 17.
@@
Case II is burdened with complex texture mapping due to variations in the texture and decoration of the garment. Consequently, an average color is padded into the target shape, resulting in blurred texture and discarded decoration. This causes an increase of 85.54% and 92.51% in ${{\rm{FID}}\left({\tilde{\cal I}_{TAB}^{hc},\tilde{\cal I}_{SAB}^{hc}}\right)}$ and ${{\rm{FID}}\left({\tilde{\cal I}_{SB}^{hc},{\cal I}_{B}^{hc}}\right)}$, respectively. As TPS deformation has a low DOF, TPS deformation is skilled in warping a coarse result for virtual try-on garments with standard poses and large scales, but it is hard to tackle the fine warping from pose to pose for garment transfer. Therefore, case III cannot accurately warp the garment to match the target shape and tends to produce distortion, leading to an increase of 175.44% and 341.51% in ${{\rm{MSE}}\left({\tilde{\cal I}_{TAB}^{hc},\tilde{\cal I}_{SAB}^{hc}}\right)}$ and ${{\rm{FID}}\left({\tilde{\cal I}_{TAB}^{hc},\tilde{\cal I}_{SAB}^{hc}}\right)}$, respectively. In case IV, rendering texture is also a challenge, and the poor warped result and unaligned ${{\cal I}_{B}^{hc}}$ cause training instability and collapse, resulting in a reduction of ${{\rm{IS}}\left({\tilde{\cal I}_{SAB}^{hc}}\right)}$ by 24.01%.
@@
In case II, position mapping and shape adjustment are all conducted by high-DOF flow warping. However, when ${{\cal M}_{B}^{hc}}$ is significantly smaller than ${{\cal M}_{SAB}^{hc}}$, it tends to excessively stretch the local region to match the target shape precisely. In contrast, case I aligns ${{\cal I}_{B}^{hc}}$ globally with ${{\cal M}_{SAB}^{hc}}$ to achieve a similar scale and position. Subsequently, flow warping is leveraged to adjust the shape accurately. As a result, the texture blur and garment style variation are reduced, and ${{\rm{MSE}}\left({\tilde{\cal I}_{TAB}^{hc},\tilde{\cal I}_{SAB}^{hc}}\right)}$ and ${{\rm{LPIPS}}\left({\tilde{\cal I}_{TAB}^{hc},\tilde{\cal I}_{SAB}^{hc}}\right)}$ decrease by 4.32% and 2.27%. In case III, the FSV modulates the pose and garment features to estimate a coarse try-on result, which performs well in virtual try-on without parsing reasoning. However, our garment transfer pipeline provides the fined target shape, rendering the mechanism of FSV unnecessary. Conversely, the innovation of case I learns correspondence at the shape and content level, emphasizing semantic alignment and consistency in garment warping. As a result, ~${\tilde{\cal I}_{SAB}^{hc}}$ is more realistic, and ${{\rm{IS}}\left({\tilde{\cal I}_{SAB}^{hc}}\right)}$ and ${{\rm{IS}}\left({\tilde{\cal I}_{SB}^{hc}}\right)}$ increase by 2.05% and 3.15%.
@@ 4.2.3 Arm Regrowth
The task of arm regrowth can be divided into two sub-tasks: propagating the original arm part and reasoning the new arm part. In case II, the augmented ${{\cal I}_{A}^{ha}}$ is misaligned with ${{\cal M}_{A}^{ha}}$ through an affine transformation, resulting in ${{G_{r}}}$ padding the overall target shape with the semantic features of ${{\cal I}_{A}^{ha}}$. Consequently, the results of case II discard high-frequency details of the hand and neck via inferring rather than content propagating at the overlapped region between ${{\cal M}_{SAB}^{ha}}$ and ${{\cal M}_{A}^{ha}}$. In case III, new arm exposure is more realistically simulated in the training phase by using random cropping, resulting in a relative reduction of 18.85% and 14.54% in MSE and LPIPS respectively when compared to case II. Nevertheless, the results of case III still exhibit artifacts and blurring. Case IV further impedes the propagation of the original arm part by using random flipping, resulting in worse results compared to case III. Unlike these data augmentation techniques, case I aims to enable the input conditions of the training and testing phase to be identical and performs well in both sub-tasks, leading to significant quantitative results.
@@ 4.3.1 Virtual Try-on Method
Baseline We compare our method with virtual try-on methods in garment transfer. Virtual try-on methods consist of VITON [1], CP-VTON [2], ACGPN [5], DCTON [8], PF-AFN [9], and FSV [11].
Metrics We compare these methods to transfer the garment from ${{{\cal I}_{B}}}$ to ${{{\cal I}_{A}}}$ to obtain ~${{\tilde{\cal I}_{AB}}}$. In order to perform this transfer, virtual try-on methods are conditioned on both ${{{\cal I}_{A}}}$ and ${{\cal I}_{B}^{hc}}$. As the ground truth for ~${{\tilde{\cal I}_{AB}}}$ is unavailable, we present several examples in Fig. 21 and evaluate their subjective image quality using IS [35] and hyperIQA [36] in Table 4. To conduct a more detailed quantitative comparison, we re-transfer the garment from ~${{\tilde{\cal I}_{AB}}}$ to ${{{\cal I}_{B}}}$ and utilize ${{{\cal I}_{B}}}$ as the ground truth. We also adopt the same metrics as in Sec. 4.2.2. Specifically, we utilize the pre-trained Grapy-ML to extract ~${{\tilde{\cal M}_{AB}}}$ from ~${{\tilde{\cal I}_{AB}}}$ and yield the transferred garment ~${\tilde{\cal I}_{AB}^{hc}}$. Subsequently, virtual try-on methods are conditioned on both ~${\tilde{\cal I}_{AB}^{hc}}$ and ${{{\cal I}_{B}}}$ to predict ~${{\tilde{\cal I}_{B}}}$, and the quantitative results are presented in Table 4.
@@
[…]
Method ${{{\cal I}_{B}}\to{{\cal I}_{A}}}$${{\tilde{\cal I}_{AB}}\to{{\cal I}_{B}}}$
IS↑ hyperIQA↑ MSE↓ PSNR↑ SSIM↑ FID↓ LPIPS↓ IS↑
Ours 2.75 35.94 389.029 23.57 0.863 19.14 0.108 2.734
VITON 2.39 31.81 3867.18 16.15 0.700 75.91 0.258 2.374
[…]
ACGPN 2.50 27.11 647.72 21.31 0.825 27.64 0.147 2.496
DCTON 2.42 29.49 770.71 20.27 0.814 32.18 0.159 2.455
PF-AFN 2.31 28.80 1359.94 18.15 0.823 46.15 0.176 2.280
FSV 2.38 27.92 1711.46 17.12 0.820 56.50 0.191 2.317
Virtual try-on methods are incapable of capturing the position and style features of transferred garments, resulting in misalignment between warped garments and transfer parsing while yielding artifacts in the unaligned region. In contrast, our method precisely warps ${{\cal I}_{B}^{hc}}$ towards ${{\cal M}_{SAB}^{hc}}$ through initial aligning and flow warping. Additionally, these virtual try-on methods cannot avoid being affected by the garment shape features of ${{{\cal I}_{A}}}$ when reasoning a proper transfer parsing, leading to ${{\cal I}_{B}^{hc}}$ warping toward ${{\cal I}_{A}^{hc}}$. In contrast, our method infers a parsing that entangles the shape feature of the transfer garment with the pose feature of another person, while maintaining their inherent features. This allows our results to be more aligned with the intended garment transfer. In quantitative comparisons, our method achieves the highest scores of IS and hyperIQA in garment transfer from ${{{\cal I}_{B}}}$ to ${{{\cal I}_{A}}}$. Moreover, our MSE and FID are approximately 20.89% and 34.37% lower than those of other virtual try-on methods in garment transfer from ~${{\tilde{\cal I}_{AB}}}$ to ${{{\cal I}_{B}}}$.